# Topological Weyl Altermagnetism in CrSb

### v1

@@ Methods
DFT calculations Full relativistic spin-polarized electronic structure and Fermi surface calculations are done using the full potential local orbital code FPLO [43, 44], on a $k$-mesh of $20\times 20\times 13$ points. We use the generalized gradient approximation to the exchange and correlation potential by Perdew, Burke, and Ernzerhof (PBE). The SS and WPs calculations are performed based on the DFT calculation from VASP [45] employing the projector augmented wave method [46]. The Brillouin zone is sampled with a $10\times 10\times 8$, with Gamma-centered $k$-point. The energy cutoff of the plane wave basis is set to 550 eV. The Hubbard term was introduced and set to be 0.8 eV in the d-orbital of Cr-atom in the DFT framework (DFT+U) in order to account for the electron-electron correlation. The Wannier Hamiltonian is symmetrized based on the maximally localized Wannier functions generated by the WANNIER90 interface [47]. To locate the WPs and calculate the chirality, Wanniertools is implemented [48].
@@ Acknowledgements
C.L. and H.L. proposed the project. C.L., M.L.H., H.L., O.T. and J.V.D.B. conceived the project. C.L. carried out the ARPES experiments with the assistance from Y.W. and W.Y.C.. M.L.H. and O.J. contributed to the band structure calculations. Z.L.L. contributed to CrSb crystal growth. C.L. contributed to software development for data analysis and analyzed the data. C.L., M.L.H., H.L., O.T. and J.V.D.B. wrote the paper. B.T., M.L., C.P. and T.K. provided the beamline support. C.L, M.L.H., Y.W., H.L., C.F., M.V., O.J., O.T. and J.V.D.B. contributed to the scientific discussions. All authors participated in commented on the paper.
@@
Figure 1: Crystal structure, WPs and calculated electronic structure of CrSb. (a) The crystal structure of CrSb with the space group $P6_{3}/mmc$ (no. 194). (b) The 3D BZ of the original unit cell of CrSb, and the corresponding two-dimensional BZ projected on the (001) plane (green lines). The red, green, blue and orange planes are the four mirror planes in the BZ on which the altermagnetic bands are spin-degenerate in absence of SOC. (c) Calculated bulk Fermi surface in the first BZ. The colormap from blue to orange of Fermi surface sheet represents the expectation value of spin along the $z$ direction. (d-e) Illustration of the two types of WPs: opposite-spin (d) and same-spin (e). Top (f) and side (g) views of the distribution of same-spin WPs in the 3D BZ. Red dots represent nodes with chirality $\chi=+1$, green dots $\chi=-1$; the spin polarizations are indicated by cyan and pink. The blue line represents spin-up and spin-down degeneracy. (h-i) The calculated SS of (001) on Sb (h) and Cr (i) terminations along $\rm\overline{M}_{1}-\overline{\Gamma}-\overline{M}_{2}-\overline{K}_{2}-\overline{\Gamma}-\overline{K}_{1}-\overline{M}_{1}$. The WPs are marked by the red ($\chi=1$) and green ($\chi=-1$) dots. Star, square and round dots represent the opposite-spin WPs on the high symmetry line/plane (WP1), opposite-spin WPs on general momenta (WP2) and same-spin WPs (WP3), respectively.
Figure 2: Fermi surface of CrSb. (a-b) projected Fermi surfaces for Sb (a) and Cr (b) surface terminations. (c-h) Calculated bulk Fermi surfaces at kz = 1 $\pi/c$ (c), 0.8 $\pi/c$ (d), 0.6 $\pi/c$ (e), 0.4 $\pi/c$ (f), 0.2 $\pi/c$ (g) and 0 $\pi/c$ (h) planes, integrated over a ±$\pm 0.1$ $\pi/c$ interval. (i) Photon energy dependent ARPES spectral intensity map at Fermi level along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$. (j) corresponding calculated kz-kx Fermi surface. (k) Photon energy dependent Fermi surfaces measured with photon energies of 97 eV (i), 102 eV (ii), 107 eV (iii), 112 eV (iv), 117 eV (v), 122 eV (vi) on an uneven surface. (l-m) The same measurements on flat areas of Sb (l) and Cr (m) terminated surfaces. The SSs are marked by orange arrows. The corresponding sample photos are shown in the bottom left corner of k(i), l(i) and m(i).

### v2

@@ Methods
DFT calculations Full relativistic spin-polarized electronic structure and Fermi surface calculations are done using the full potential local orbital code FPLO [51, 52], on a $k$-mesh of $20\times 20\times 13$ points. We use the generalized gradient approximation to the exchange and correlation potential by Perdew, Burke, and Ernzerhof (PBE). The surface state and WPs calculations are performed based on the DFT calculation from VASP [53] employing the projector augmented wave method [54]. The Brillouin zone is sampled with a $10\times 10\times 8$, with Gamma-centered $k$-point. The energy cutoff of the plane wave basis is set to 550 eV. The Hubbard term was introduced and set to be 0.8 eV in the $d$ orbitals of Cr-atom in the DFT framework (DFT+U) in order to account for the electron-electron correlation. The Wannier based Hamiltonian is symmetrized based on the maximally localized Wannier functions generated by the WANNIER90 interface [55]. The projectors are $d$ orbitals of Cr and $p$ orbitals of Sb atoms with the well-fitted region from -2 to 2 eV. To locate the WPs and calculate the chirality, Wanniertools is implemented [56].
@@ Acknowledgements
C.L. and H.L. proposed the project. C.L., M.L.H., H.L., O.T. and J.V.D.B. conceived the project. C.L. carried out the ARPES experiments with the assistance from Y.W. and W.Y.C.. M.L.H. and O.J. contributed to the band structure calculations. Z.L.L. contributed to CrSb crystal growth. C.L. contributed to software development for data analysis and analyzed the data. C.L., M.L.H., H.L., O.T. and J.V.D.B. wrote the paper. B.T., M.L., C.P. and T.K. provided the beamline support. C.L, M.L.H., Y.W., H.L., C.F., M.V., O.J., O.T. and J.V.D.B. contributed to the scientific discussions. All authors participated in and commented on the paper.
@@
Figure 1: Crystal structure, WPs and calculated electronic structure of CrSb. (a) The crystal structure of CrSb with the space group $P6_{3}/mmc$ (no. 194). (b) The 3D BZ of the original unit cell of CrSb, and the corresponding two-dimensional BZ projected on the (001) plane (green lines). The red, green, blue and orange planes are the four mirror planes in the BZ on which the altermagnetic bands are spin-degenerate in absence of SOC. (c) Calculated bulk Fermi surface in the first BZ. The colormap from blue to orange of Fermi surface sheet represents the expectation value of spin along the $z$ direction. (d-e) Illustration of the two types of WPs: opposite-spin (d) and same-spin (e). Top (f) and side (g) views of the distribution of same-spin WPs in the 3D BZ. Red dots represent nodes with chirality $\chi=+1$, green dots $\chi=-1$; the spin polarizations are indicated by cyan and pink. The blue line represents spin-up and spin-down degeneracy. (h-i) The calculated surface state of (001) on Sb (h) and Cr (i) terminations along $\rm\overline{M}_{1}-\overline{\Gamma}-\overline{M}_{2}-\overline{K}_{2}-\overline{\Gamma}-\overline{K}_{1}-\overline{M}_{1}$. The WPs are marked by the red ($\chi=1$) and green ($\chi=-1$) dots. Star, square and round dots represent the opposite-spin WPs on the high symmetry line/plane (WP1), opposite-spin WPs on general momenta (WP2) and same-spin WPs (WP3), respectively.
Figure 2: Fermi surface of CrSb. (a-b) projected Fermi surfaces for Sb (a) and Cr (b) surface terminations with WPs overlapped. Red dots represent nodes with chirality $\chi=+1$, green dots $\chi=-1$; (c-h) Calculated bulk Fermi surfaces at kz = 1 $\pi/c$ (c), 0.8 $\pi/c$ (d), 0.6 $\pi/c$ (e), 0.4 $\pi/c$ (f), 0.2 $\pi/c$ (g) and 0 $\pi/c$ (h) planes, integrated over a ±$\pm 0.1$ $\pi/c$ interval. (i) Photon energy dependent ARPES spectral intensity map at Fermi level along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$. (j) corresponding calculated kz-kx Fermi surface. (k) Photon energy dependent Fermi surfaces measured with photon energies of 97 eV (i), 102 eV (ii), 107 eV (iii), 112 eV (iv), 117 eV (v), 122 eV (vi) on an uneven surface. (l-m) The same measurements on flat areas of Sb (l) and Cr (m) terminated surfaces. The SFAs are marked by orange arrows. The corresponding sample photos are shown in the bottom left corner of k(i), l(i) and m(i).